# Incremental Learning for Robot Shared Autonomy

### v3

@@ Abstract
Shared autonomy holds promise for improving the usability and accessibility of assistive robotic arms, but current methods often rely on costly expert demonstrations and remain static after pretraining, limiting their ability to handle real-world variations. Even with extensive training data, unforeseen challenges—especially those that fundamentally alter task dynamics, such as unexpected obstacles or spatial constraints—can cause assistive policies to break down, leading to ineffective or unreliable assistance. To address this, we propose ILSA, an Incrementally Learned Shared Autonomy framework that continuously refines its assistive policy through user interactions, adapting to real-world challenges beyond the scope of pre-collected data. At the core of ILSA is a structured fine-tuning mechanism that enables continual improvement with each interaction by effectively integrating limited new interaction data while preserving prior knowledge, ensuring a balance between adaptation and generalization. A user study with 20 participants demonstrates ILSA’s effectiveness, showing faster task completion and improved user experience compared to static alternatives. Code and videos are available at https://ilsa-robo.github.io/.
@@ I INTRODUCTION
Assistive devices such as wheelchair-mounted robotic arms have the potential to benefit millions of individuals with physical disabilities and enhance their independence [5, 2]. Shared autonomy, where human operators and autonomous systems share decision-making, has emerged as a promising way to make these devices more intuitive and effective [27, 10].
@@
To address the inability of static policies to adapt to unexpected real-world challenges that alter task execution, we introduce an Incrementally Learned Shared Autonomy (ILSA) framework (Fig. 2) that continuously fine-tunes its assistive policy as the user interacts with the system. A key technical contribution of ILSA is its structured fine-tuning mechanism, which updates the assistive policy after each user interaction, ensuring enhanced assistance in every subsequent use. This frequent adaptation is particularly challenging due to limited deployment data, which can affect model stability and generalizability. To address this, our fine-tuning mechanism consists of three key components: (1) Corrected Trajectory Supervision, which prevents the model from reinforcing suboptimal actions by generating corrected trajectories from user interactions and using them for fine-tuning; (2) Layered Supervision, which integrates new interaction data with pretraining knowledge to balance adaptation and retention, without requiring manual data reweighting; and (3) Partial Model Update, which selectively fine-tunes critical components to enhance adaptation while maintaining generalization and robustness. These designs collectively enable continual adaptation while preserving previously acquired knowledge, ensuring ILSA retains the generalization ability gained from pretraining. The effectiveness of our fine-tuning approach is validated through a quantitative ablation study.
@@
  • We design a structured fine-tuning mechanism that enables ILSA to improve its assistive policy after each user interaction, ensuring continual improvement while balancing adaptation and generalization ability. We validate its effectiveness through a quantitative ablation study.
@@ II-A Shared Autonomy with Continual Adaptation
In recent years, a limited number of shared autonomy works explored frameworks that refine assistive policies through repeated user interactions. Two notable works in this direction are Confidence-Aware Shared Autonomy (CASA) [32] and Shared Autonomy across Repeated Interactions (SARI) [12]. Both methods aim to improve assistive policies over time by expanding the robot’s skill set through learning new tasks or goals from user demonstrations. Specifically, CASA estimates confidence in a user’s intent to determine whether a task is known or novel. If recognized, it blends user and robot control; otherwise, it relinquishes control entirely, allowing the user to demonstrate the task. CASA then applies inverse reinforcement learning (IRL) to infer a cost function and incorporate the new task into its skill set. Similarly, SARI automates familiar tasks based on prior interactions while deferring control to the user for novel ones. It periodically retrains its policy on collected trajectories, enabling continual refinement over time.
@@
Furthermore, CASA and SARI require multiple user demonstrations before updating and retrain periodically on aggregated interaction data, which can lead to long periods of suboptimal performance before policy updates, limiting the user experience. In contrast, we aim to fine-tune the assistive policy after each interaction, enabling continuous adaptation in real-time. Achieving this, however, presents key challenges including ensuring stability and generalization when updating with limited new data. To address this, we focus on effectively integrating new interaction data while preserving pretraining knowledge, striking a balance between adaptation and retention. While prior shared autonomy frameworks have not explicitly explored this direction, we introduce a structured fine-tuning mechanism specifically designed for this challenge. Through our key designs, ILSA ensures that each update contributes to improved assistance, effectively enabling adaptation to real-world challenges beyond the scope of pre-collected data with minimal interaction overhead.
@@ II-B Human-in-the-loop Machine Learning
Another related area is human-in-the-loop machine learning [8, 1], where autonomous agents refine their policies through human interactions. A variety of human feedback modalities—such as demonstrations, preferences, rankings[4], interventions [14, 21, 28], and scalar-valued critiques [20, 29]—have been explored to enhance autonomous decision-making. Among these approaches, interactive imitation learning [6, 24, 3], which incorporates human interventions into imitation learning, is particularly relevant to our work. A notable example is DAgger (Dataset Aggregation)[26] and its variants, such as EnsembleDAgger[23] and HG-DAgger [14], which address covariate shift in imitation learning by iteratively collecting expert interventions on novel states encountered during deployment, improving policy robustness.
@@ III-B Simulated Kinematic Trajectories Generation
To further simulate user interaction, we generate synthetic user actions $a_{u,t}$ to mimic user inputs. The user action is defined as a 3D vector representing intended end-effector movements in the $x$, $y$, and $z$ directions. Each dimension corresponds to a normalized velocity command: if the target position in dimension $i$ is greater than the current position, we sample $a_{u,t,i}$ from $[0,1]$; otherwise, we sample from $[-1,0]$. In real-world deployment, these inputs are provided by a user via a dual-joystick Xbox controller, where one joystick operates two translational axes (e.g., $x$ and $y$), and the other controls the third axis ($z$).
@@ III-C Robot Action Generation Model
To pretrain $\mathcal{F}_{\theta}$, we use a set of simulated kinematic trajectories (Section III-B) and optimize a loss function designed to balance task completion and user intent alignment. This function includes terms ($\mathcal{L}_{demo,m}$, $\mathcal{L}_{demo,f}$) that align both intermediate and final actions with training data to ensure task completion, along with additional constraints ($\mathcal{L}_{direc}$, $\mathcal{L}_{order}$) to maintain consistency with user inputs. Since this model serves as the basis for ILSA’s assistive policy rather than its core technical contribution, which focuses on incremental adaptation to real-world challenges, we defer the full mathematical formulation of the loss functions and implementation details to Appendix A.
@@ III-D1 Corrected Trajectory Supervision
As illustrated in this example, raw trajectories contain both suboptimal policy actions and user corrections that wouldn’t be necessary without the suboptimal actions. Directly using these trajectories for training risks misleading the model into treating these corrective actions as optimal behavior to be replicated.
To address this, we generate a “corrected trajectory” for each raw trajectory observed during human use. As shown in Fig. 3, when the user overwrites $a_{f,t}$ (red segments in Fig. 3), let $t_{1}$ be the start of the overwrite and $t_{2}$ the time when the robot resumes executing $a_{f,t}$. Before $t_{1}$, if the robot executes $a_{f,t}$ for a period (blue segments), let $t_{0}$ be the start of that period. Let $p_{t_{0}}$, $p_{t_{1}}$, and $p_{t_{2}}$ represent the robot’s positions at these times. An example of these key positions is marked in Fig. 3.
@@ III-D2 Layered Supervision
To ensure a smooth user experience, ILSA updates its policy after each interaction. However, frequent updates mean that each training step relies on only a small set of newly collected data, making the system susceptible to overfitting and catastrophic forgetting of previously acquired knowledge [9, 22, 15].
@@
    Final Output Supervision: The new user-generated trajectories are used exclusively to supervise the final output action $a_{f,t}$, ensuring adaptation to real-world interactions.
@@
This approach, illustrated by the red arrows in Fig. 2, ensures that ILSA remains adaptive while preserving knowledge from pretraining. Compared to explicit weighting schemes, layered supervision provides a more principled way to balance adaptation and generalization.
@@ IV-A Tasks and Experiment Settings
All experiments are conducted with a 6-DOF Kinova Gen3 robotic arm, and users use an Xbox controller to send commands to the robot.
@@ IV-A2 Pill Bottle Storage
In this task, the spatial constraints between the drawer and the pill bottle were not accounted for in pretraining, requiring ILSA to adapt its motion strategy to avoid unintended collisions.
@@ IV-B Ablation Study
For the ablation study, one researcher conducted five experiments for each model variant on the cereal pouring task. The tested designs were randomized and interleaved across experiments. To evaluate adaptation across diverse task layouts, object positions were randomized for each trial, resulting in a total of 20 unique task layouts across the five experiments. For fair comparison, all model variants were evaluated on the same sequence of object positions within each experiment. This ensures that performance differences arise from model adaptation rather than variations in task difficulty.
Across both the ablation and user studies (Section II-B), we primarily  [^1] Additional experimental results are available on our project website. evaluate task completion time as a direct and interpretable measure of user efficiency and system adaptability, which is particularly relevant in assistive settings where efficiency and ease of use directly impact user experience.
@@ V CONCLUSIONS
Appendix A Loss Function Details

### v4

@@ Abstract
Shared autonomy holds promise for improving the usability and accessibility of assistive robotic arms, but current methods often rely on costly expert demonstrations and remain static after pretraining, limiting their ability to handle real-world variations. Even with extensive training data, unforeseen challenges—especially those that fundamentally alter task dynamics, such as unexpected obstacles or spatial constraints—can cause assistive policies to break down, leading to ineffective or unreliable assistance. To address this, we propose ILSA, an Incrementally Learned Shared Autonomy framework that continuously refines its assistive policy through user interactions, adapting to real-world challenges beyond the scope of pre-collected data. At the core of ILSA is a structured fine-tuning mechanism that enables continual improvement with each interaction by effectively integrating limited new interaction data while helping to preserve prior knowledge, aiming for a balance between adaptation and generalization. A user study with 20 participants demonstrates ILSA’s effectiveness, showing faster task completion and improved user experience compared to static alternatives. Code and videos are available at https://ilsa-robo.github.io/.
@@ I INTRODUCTION
Assistive devices such as wheelchair-mounted robotic arms have the potential to benefit millions of individuals with disabilities and enhance their independence [3, 2]. Shared autonomy, where human operators and autonomous systems share decision-making, has emerged as a promising way to make these devices more intuitive and effective [22, 6].
@@
To address the inability of static policies to adapt to unexpected real-world challenges that alter task execution, we introduce an Incrementally Learned Shared Autonomy (ILSA) framework (Fig. 2) that continuously fine-tunes its assistive policy as the user interacts with the system. A key technical contribution of ILSA is its structured fine-tuning mechanism, which updates the assistive policy after each user interaction, ensuring enhanced assistance in every subsequent use. This frequent adaptation is particularly challenging due to limited deployment data, which can affect model stability and generalizability. To address this, our fine-tuning mechanism consists of three key components: (1) Corrected Trajectory Supervision, which prevents the model from reinforcing suboptimal actions by generating corrected trajectories from user interactions and using them for fine-tuning; (2) Layered Supervision, which integrates new interaction data with pretraining knowledge to balance adaptation and retention, without requiring manual data reweighting; and (3) Partial Model Update, which selectively fine-tunes critical components to enhance adaptation while maintaining generalization and robustness. These designs collectively enable continual adaptation while helping to preserve previously acquired knowledge, thus supporting ILSA’s generalization ability gained from pretraining. Our fine-tuning approach is validated through a quantitative ablation study.
@@
  • We design a structured fine-tuning mechanism that enables ILSA to improve its assistive policy after each user interaction, helping to achieve continual improvement while balancing adaptation and generalization ability. We validate its effectiveness through a quantitative ablation study.
@@ II-A Shared Autonomy with Continual Adaptation
In recent years, a limited number of shared autonomy works explored frameworks that refine assistive policies through repeated user interactions. Two notable works in this direction are Confidence-Aware Shared Autonomy (CASA) [25] and Shared Autonomy across Repeated Interactions (SARI) [8]. Both methods aim to improve assistive policies over time by expanding the robot’s skill set through learning new tasks or goals from user demonstrations. Specifically, CASA estimates confidence in a user’s intent to determine whether a task is known or novel. If a task is unrecognized, it relinquishes control entirely, allowing the user to demonstrate the task. CASA then applies inverse reinforcement learning (IRL) and incorporate the new task into its skill set. Similarly, SARI automates familiar tasks based on prior interactions while deferring control to the user for novel ones. It periodically retrains its policy on collected trajectories, enabling continual refinement over time.
@@
Furthermore, CASA and SARI require multiple user demonstrations before updating and retrain periodically on aggregated interaction data, which can lead to long periods of suboptimal performance before policy updates, limiting the user experience. In contrast, we aim to fine-tune the assistive policy after each interaction, enabling continuous adaptation in real-time. To achieve this, we focus on effectively integrating new interaction data while preserving pretraining knowledge, striking a balance between adaptation and retention. While prior shared autonomy frameworks have not explicitly explored this direction, we introduce a structured fine-tuning mechanism specifically designed for this challenge. Through our key designs, ILSA ensures that each update contributes to improved assistance, effectively enabling adaptation to real-world challenges beyond the scope of pre-collected data with minimal interaction overhead.
@@ II-B Human-in-the-loop Machine Learning
Another related area is human-in-the-loop machine learning [5, 1], where autonomous agents refine their policies through human interactions. A notable example is DAgger (Dataset Aggregation)[21] and its variants, such as EnsembleDAgger[19] and HG-DAgger [10], which address covariate shift in imitation learning by iteratively collecting expert interventions on novel states encountered during deployment, improving policy robustness.
@@ III-B Simulated Kinematic Trajectories Generation
To further simulate user interaction, we generate synthetic user actions $a_{u,t}$ to mimic user inputs. The user action is defined as a 3D vector representing intended end-effector movements in the $x$, $y$, and $z$ directions. Each dimension corresponds to a normalized velocity command: if the target position in dimension $i$ is greater than the current position, we sample $a_{u,t,i}$ from [0,1][0,1][ 0 , 1 ]; otherwise, we sample from [1,0][-1,0][ - 1 , 0 ]. In real-world deployment, these inputs are provided by a user via a dual-joystick Xbox controller, where one joystick operates two translational axes (e.g., $x$ and $y$), and the other controls the third axis ($z$). Note that for usability, in most cases, users control only translation (3D) while the robot autonomously generates the end-effector orientation. When direct rotational control is necessary, the user’s input bypasses the policy and is executed directly. This design reduces user burden and was validated in pilot studies.
@@ III-C Robot Action Generation Model
To pretrain $\mathcal{F}_{\theta}$, we use a set of simulated kinematic trajectories (Section III-B) and optimize a loss function designed to balance task completion and user intent alignment. This function includes terms ($\mathcal{L}_{demo,m}$, $\mathcal{L}_{demo,f}$) that align both intermediate and final actions with training data to ensure task completion, along with additional constraints ($\mathcal{L}_{direc}$, $\mathcal{L}_{order}$) to maintain consistency with user inputs. Since this model serves as the basis for ILSA’s assistive policy rather than its core technical contribution, which focuses on incremental adaptation to real-world challenges, we defer the full implementation details—including network architecture and training configuration—to Appendix A.
@@ III-D1 Corrected Trajectory Supervision
As illustrated in this example, raw trajectories contain both suboptimal policy actions and user corrections that wouldn’t be necessary without the suboptimal actions. Directly using these trajectories for training risks misleading the model into treating these corrective actions as optimal behavior.
To address this, we generate a “corrected trajectory” for each raw trajectory observed during human use. As shown in Fig. 3, when the user overwrites $a_{f,t}$ (red segments in Fig. 3), let $t_{1}$ be the start of the overwrite and $t_{2}$ the time when the robot resumes executing $a_{f,t}$. Before $t_{1}$, if the robot executes $a_{f,t}$ for a period (blue segments), let $t_{0}$ be the start of that period. Let $p_{t_{0}}$, $p_{t_{1}}$, and $p_{t_{2}}$ represent the robot’s positions at these times.
@@ III-D2 Layered Supervision
To ensure a smooth user experience, ILSA updates its policy after each interaction. However, frequent updates mean that each training step relies on only a small set of newly collected data, making the system susceptible to overfitting and catastrophic forgetting of previously acquired knowledge [18, 11].
@@
    Final Output Supervision: The new user-generated trajectories are used exclusively to supervise the final output action $a_{f,t}$, ensuring adaptation to new data.
@@
This approach, illustrated by the red arrows in Fig. 2, helps ILSA remain adaptive while mitigating the risk of catastrophic forgetting and preserving knowledge from pretraining. Compared to explicit weighting schemes, layered supervision provides a more principled way to balance adaptation and generalization.
@@ III-D3 Partial Model Update
In each incremental update, we fine-tune the model for 10 epochs on the new and pretraining data, with this value empirically set for stable adaptation.
@@ IV-A Tasks and Experiment Settings
All experiments are conducted with a 6-DOF Kinova Gen3 robotic arm.
@@ IV-A2 Pill Bottle Storage
In this task, the spatial constraints between the drawer and the pill bottle were not accounted for in pretraining, requiring ILSA to adapt its motion strategy to avoid collisions.
@@ IV-B Ablation Study
For the ablation study, one researcher conducted five experiments for each model variant on the cereal pouring task. The tested designs were randomized and interleaved across experiments. Object positions were randomized for each trial, resulting in a total of 20 unique task layouts across the five experiments. All model variants were evaluated on the same sequence of object positions within each experiment. This ensures that performance differences arise from model adaptation rather than variations in task difficulty. For all methods, after each trial, we fine-tuned for a fixed number of epochs, with the value empirically set for each variant to ensure stable adaptation.
Across both the ablation and user studies (Section II-B), we primarily  [^1] Additional experimental results are available on our project website. evaluate task completion time as a direct measure of user efficiency and system adaptability.
@@ V CONCLUSIONS
NIST Disclaimer
Certain commercial entities, equipment, or materials may be identified in this document in order to illustrate a point or concept. Such identification is not intended to imply recommendation or endorsement by NIST, nor is it intended to imply that the entities, materials, or equipment are necessarily the best available for the purpose.
Appendix A Implementation Details
A-A Network Architecture
ILSA’s robot action generation model consists of a transformer encoder and three multilayer perceptrons (MLPs). Specifically, we adopt the transformer structure implemented in the Robomimic framework [17]. The three MLPs each have two hidden layers with 1024 units per layer, followed by ReLU activations. All model components are trained jointly using the loss functions described below.
A-B Loss Functions